# CAAP: Context-Aware Action Planning Prompting to Solve Computer Tasks with Front-End UI Only

### v1

@@ Appendix A Train datasets
Data collection for fine-tuning the Visual Observer involves a two-step process: screenshot collection and annotation of the collected UI elements. Initially, the target task is manually performed while screenshots are taken at each step. Image comparisons are then conducted to remove duplicate screenshots. From 10 episodes for each of the 67 tasks, a total of 1,768 screenshots were collected. In the second stage, a GUI-based annotation tool, as depicted in Figure 7, is utilized. Each UI element within the screenshots is marked with bounding boxes, and specific attributes are assigned. These attributes are subsequently stored in a JSON format. Data augmentation techniques are also employed to enhance the dataset. These techniques include color adjustments, minor scaling of bounding box sizes, and the random insertion of Gaussian noise. As a result, the total number of annotated data items is tripled compared to the original dataset.
@@
We generated between 0 and 5 demonstrations per task, tailored to the task’s complexity, amassing a total of 99 demonstrations for our experiments. We recorded the trajectory of each action performed by the demonstrators, alongside screenshots taken at the time of each action. These screenshots were subsequently analyzed to extract UI elements, enabling the identification of the UI component interacted with during mouse-related actions, such as clicks.
@@
Appendix B Element coverage of CAAP agent
@@
Table 2 provides a comprehensive overview of the types of frontend elements handled by the agent, predominantly including well-known types of frontend objects. Beyond the traditional categories, this study introduces the concept of ’tabled text’ to describe texts commonly embedded within table structures, and presents ’draggable text’ to represent text elements of which its text can be partially selected. These additions ensure extensive coverage of both interactive and static elements found in varying types of applications, thereby enhancing the understanding capability of the visual environmental states. Additionally, our agent covers conventional shapes and icons. The shape type includes circle, triangle, and rectangle, distinguished by their subtype values. Moreover, the icon type encompasses subtypes such as "back," "delete," "important," "forward," "reply," "retweet," "like," "share," "menu," "search," and "send," as identified in MiniWob++.
@@
Appendix C Action coverage of CAAP agent
Tables 3 and Table 4 detail the action types supported by the CAAP agent and the requisite parameters, respectively. The action executer operates using atom actions based on mouse and keyboard interactions, specifically: mouse down, mouse move, mouse up, and type a key. However, when conveying candidate actions to an LLM, it is more effective to transmit them as high-level actions capable of altering the environment. Consequently, as indicated in Table 3, actions have been delineated into units such as clicking on an element, clicking an empty position, clicking on an element during pressing control key, and typing text.
@@ Appendix C Action coverage of CAAP agent
Appendix D Case study
@@
Appendix E Annotation tool
@@
Appendix F Prompt examples
@@
Figure 8: An prompt example of generating a rationale for an agent action.
@@
{mdframed}
[leftmargin=5pt,rightmargin=5pt]
### The UI Element List of the Current Screen That You Can Take Next Actions On ###
(Note: Coordinates are given in the form: center_x [left_edge_x-right_edge_x], center_y [top_edge_y-bottm_edge_y])
- ID: element_1, data: {type: dropdown, X: 76 [1-152], Y: 66 [55-76], text: "Theodora", focused: False}
- ID: element_2, data: {type: button, X: 50 [1-98], Y: 96 [80-112], text: "Submit", focused: False}
### Instructions ###
What should be the next actions(action_2, action_3, …) that can be performed on the current screen?
If there is an action that can complete the task, perform it immediately. It’s better if you can complete the task with fewer actions.
Your answer must be composed of the following five sections:
First, explain in detail what has been done so far up to action_1 and analyze why these steps were needed.
Do not assume that the user could have made a mistake.
Secondly, describe every single screen component that contains information that helps user solve the task or that needs to be interacted with. Explain about the components step by step in a detailed manner considering the given task.
When the task deals with a list of items (e.g. finding the size of a group, identifying N-th item in order, etc.), you must include the full iteration of each and every items, like this: (1)first_item, (2)second_item, …, (N)N-th_item.
Thirdly, describe what needs to be done to complete the task, detailing each action from start to finish. Then, identify which steps can currently be performed based on the UI elements visible on the screen, and mention the ID of these elements.
If there is a demo available, first describe the sequence of actions demonstrated, then link these actions to the steps in your plan before outlining the complete action plan.
Lastly, each action must be in the form of Action_2=(Action: functions.some_function_name, Argument: {property_name: property_val}).
Return the actions that need to be performed on the current screen.
The actions must be separated by new line characters.
In case there are three actions to be performed, your response will be in the following form:
Action_2=(Action: functions.some_function_name, Argument: {property_name: property_val})
Action_3=(Action: functions.some_function_name, Argument: {property_name: property_val})
Action_4=(Action: functions.some_function_name, Argument: {property_name: property_val})
- MAKE SURE that when you recommend actions that need to interact with UI elements, the UI elements MUST BE in the current screen.
- MAKE SURE that when satisfying the task completion conditions, click the ’Submit’ button to finish.
- MAKE SURE that even if you do not find a suitable action, return the most plausible one.
- MAKE SURE that since texts in elements are extracted by recognizing with OCR, solve the given task considering that some mis-typos could exist.
Figure 9: An example of a CAAP prompt. Since this example demonstrates a prompt for the first LLM request to solve the given task, the last LLM response, depicted in Figure 3, is not included. The response text of this request should be included in the location between the ’Task Description’ and ’Action History’ of the next prompt as ’Last LLM Response.’
{mdframed}
[leftmargin=5pt,rightmargin=5pt]
### Analysis of Action History
So far, the only action taken has been the initiation of the task (action_1: {name: start}). This initial action is necessary to begin the process of selecting "Helli" from the dropdown list and submitting the choice. No specific interactions with the UI elements have been made yet.
### Description of Screen Components
1. **Dropdown Menu (element_1)**: This element is a dropdown menu currently displaying "Theodora" as the selected item. Its coordinates are given as X: 76 [1-152], Y: 66 [55-76], indicating its position on the screen. The dropdown is not focused, meaning it hasn’t been interacted with yet. This element is crucial for the task as it contains the list of names from which "Helli" needs to be selected.
2. **Submit Button (element_2)**: This is a button with the text "Submit" located at X: 50 [1-98], Y: 96 [80-112]. This button is essential for completing the task as it needs to be clicked after selecting "Helli" from the dropdown menu to confirm the selection.
### Completion Plan
To complete the task, the following steps need to be taken:
1. Click on the dropdown menu to expand it and reveal the list of names.
2. Select "Helli" from the list of names in the dropdown.
3. Click the "Submit" button to confirm the selection.
Based on the current screen, the actions that can be performed are:
- Clicking on the dropdown menu (element_1) to expand it and reveal the list of names, including "Helli".
- After selecting "Helli", clicking on the "Submit" button (element_2) will be necessary, but this can only be done after "Helli" has been successfully selected.
Given the current screen elements, the next action is to interact with the dropdown menu to expand it and reveal the list of names.
### Actions to be Performed
Action_2=(Action: functions.click_element, Argument: {element_id: 1})
@@
Appendix G Function calling example
@@
Appendix H Simulation result
H.1 Task Coverage
Table 5: Comparison on the number of stably solved MiniWoB++ tasks and on the overall success rate. The standard set indicates the task set covered by $C$-Net.
Method Experimented Tasks over Tasks over Tasks over Success rate
tasks 70% success 80% success 90% success on the standard set
CAAP (Ours) 67 65 61 55 63.3%
RCI 54 49 48 48 50.8%
AdaPlanner 53 51 47 36 49.2%
WebGUM 57 51 50 50 52.7%
$C$-Net 100 96 91 83 94.5%
Pix2Act 58 56 55 52 55.8%
SeeClick 55 33 24 18 38.2%
Human 100 100 98 83 94.6%
H.2 Task-level result
H.2.1 Per-task results for ablation study on CAAP components
Table 6: Per-task success rates with respect to varying combinations of prompt components.
@@ H.2.1 Per-task results for ablation study on CAAP components
Task Full Demo without CoT phrases CoT phrases with No CoT,
configurations CoT phrases without demo action-only demo no demo
@@
Mean success rate 0.944 0.920 0.857 0.911 0.809
H.2.2 Comparison of the CAAP and RCI promptings
Table 7: Per-task success rates for RCI agent with CAAP and original RCI agent.
Task RCI agent with CAAP Original RCI agent
@@ H.2.2 Comparison of the CAAP and RCI promptings
Task RCI agent with CAAP Original RCI agent
@@
Mean success rate 0.840 0.769
H.2.3 Comparison of the CAAP agent with SoTAs in per-task result
Table 8: Comparison with other works in terms of success rate per task for MiniWoB++ problems.
@@ H.2.3 Comparison of the CAAP agent with SoTAs in per-task result
Task CAAP SeeClick Pix2Act $C$-Net WebGUM AdaPlanner RCI Human
@@
Task CAAP SeeClick Pix2Act $C$-Net WebGUM AdaPlanner RCI Human
@@
Mean success rate 0.944 0.694 0.962 0.936 0.925 0.929 0.940 0.935

### v2

@@ Appendix A MiniWoB++ Experiment Details
Data collection for fine-tuning the visual observer involves a two-step process: screenshot collection and annotation of the collected UI elements. Initially, the target task is manually performed while screenshots are taken at each step. Image comparisons are then conducted to remove duplicate screenshots. From 10 episodes for each of the 73 tasks, a total of 1,768 screenshots were collected. In the second stage, a GUI-based annotation tool, as depicted in Figure 6, is utilized. Each UI element within the screenshots is marked with bounding boxes, and specific attributes are assigned. These attributes are subsequently stored in a JSON format. Data augmentation techniques are also employed to enhance the dataset. These techniques include color adjustments, minor scaling of bounding box sizes, and the random insertion of Gaussian noise. As a result, the total number of annotated data items is tripled compared to the original dataset.
@@
Train and test episode split
To differentiate between the train datasets and the test dataset, we assign distinct seed numbers when generating individual task instances within MiniWoB++. Image annotation data was collected from tasks with seed numbers ranging from 1000 to 2999, while demonstration data was sourced from seeds ranging from 3000 to 3999. Testing was conducted using seeds ranging from 0 to 999.
Appendix B Comparisons to other approaches
Table 4 summarizes the input types, agent architectures, training methodologies, and whether experimental results on MiniWoB++ and WebShop are reported for our method and other approaches.
Table 4: Comparison of agent configurations in input types, architectures, training methodologies.
Method Modality Model Training Experiment
Vision Decision Scheme MiniWoB++ WebShop
Human - - - - -
[…]
WebGUM HTML+Image ViT-B16 Flan-T5-XL SL $\surd square-root\surd$
[…]
- / Image only - -
Pix2Act Image Pix2Struct Transformer SL+RL $\surd square-root\surd$
[…]
CAAP Image
YOLOv8,
Pix2Struct
gpt-4-0125 SL $\surd square-root\surd$
B.1 Experiment result for MiniWoB++
The data requirements and reported performance in literature are described in Table 5. One notable approach is CC-Net, which reports results across the broadest range of MiniWoB++ tasks. However, this required the collection of 2.4 million demonstrations, demanding approximately 6,300 hours of human labor, and faces challenges in real-world applications due to its high reliance on DOM information. RCI (Kim et al., 2023) and AdaPlanner (Sun et al., 2023) achieved average task SRs above 92% with fewer than 100 demonstrations, but their dependence on HTML inputs limits their task coverage. When recalculating SRs, including unsupported or unreported tasks, their performance drops to 50.8% and 49.2%, respectively. WebGUM (Furuta et al., 2023) expanded task coverage by using both HTML and image inputs. However, despite utilizing 401K demonstrations, its results only showed a 1.9% improvement over RCI, achieving an SR of 52.7%. Pix2Act (Shaw et al., 2023) was the first method to achieve high task SRs on MiniWoB++ using image inputs exclusively. It leveraged 1.3 million demonstrations, achieving over 96% success on 58 tasks, with an SR of 55.8% across the same set of 100 tasks. In contrast, SeeClick, which also used image inputs but on a smaller demonstration dataset, failed to sufficiently demonstrate problem-solving ability. Our approach represents the most advanced image-only agent, solving 73 tasks with a high SR of 94.5%, using fewer than 1.8K demonstrations. Furthermore, it achieved a 63.3% SR on the 100-task range, 7.5% higher than Pix2Act, with 61 tasks demonstrating an SR of over 80%—five more tasks than Pix2Act. (See Table 12 of Appendix J.3.) This represents a significant improvement achieved with fewer than 100 human demonstrations, resulting from the powerful reasoning capabilities of the LLM and our carefully modularized agent design.
Table 5: Comparison with other methods that solve MiniWoB++ problems. While the primary configuration of $C$-Net utilizes the DOM+Image modality, we have also included the performance of the image-only modality in this table to provide a more in-depth comparison.
Method Modality Image datasize Demo datasize Reported SR Reported tasks
Human - - - 93.5% 104
[…]
WebN-T5 HTML 0 12K 48.4% 56
[…]
WebGUM HTML+Image 0 401K 94.2% 56
$C$-Net DOM+Image 0 2.4M 93.6% 104
$C$-Net Image 0 2.4M 24.0% 104
Pix2Act Image 0 1.3M 96.2% 59
SeeClick Image 0 2.8K 69.4% 55
CAAP Image 1.76K 0.1K 94.5% 73
B.2 Experiment result for WebShop
Table 6 compares our approach with studies reporting performance on the WebShop benchmark. WebGUM was trained using both HTML and image inputs with 1K human demonstrations, achieving an average reward of 66.6, surpassing the performance of the initial model introduced with the WebShop benchmark, which was trained using an SL+RL approach. Pix2Act, designed to operate solely on image inputs, reported an average reward of 46.7 when trained with 1K human demonstrations using a combination of SL and RL techniques. Our agent, utilizing a vision model training dataset collected through an automated process and just a single human demonstration, achieved a significantly higher average reward of 62.3, outperforming Pix2Act. This result is comparable to methods that leverage both HTML and image inputs, demonstrating that our approach can effectively function across web and non-web environments with minimal human effort in dataset collection.
Table 6: Comparison with other methods that solve WebShop problems.
Method Modality Image datasize Demo datasize Task score SR
Human - - - 82.1 59.6%
WebShop HTML+Image 0 10.6K 62.4 28.7%
WebGUM HTML+Image 0 1K 67.5 45.0%
Pix2Act Image 0 1K 46.7 N/A
CAAP Image 0.3K 1 62.3 34.7%
Appendix C Element coverage of CAAP agent
@@
Table 7 provides a comprehensive overview of the types of front-end elements handled by the agent, predominantly including well-known types of front-end objects. Beyond the traditional categories, this study introduces the concept of ’tabled text’ to describe texts commonly embedded within table structures, and presents ’draggable text’ to represent text elements of which its text can be partially selected. These additions ensure extensive coverage of both interactive and static elements found in varying types of applications, thereby enhancing the understanding capability of the visual environmental states. Additionally, our agent covers conventional shapes and icons. The shape type includes circle, triangle, and rectangle, distinguished by their subtype values. Moreover, the icon type encompasses subtypes such as "back," "delete," "important," "forward," "reply," "search," and "send," as identified in MiniWob++.
@@
[…]
Appendix D Action coverage of CAAP agent
Tables 8 and Table 9 detail the action types supported by the CAAP agent and the requisite parameters, respectively. The action executor operates using atom actions based on mouse and keyboard interactions, specifically: mouse down, mouse move, mouse up, and type a key. However, when conveying candidate actions to an LLM, it is more effective to transmit them as high-level actions capable of altering the environment. Consequently, as indicated in Table 8, actions have been delineated into units such as clicking on an element, clicking an empty position, clicking on an element during pressing control key, and typing text.
@@ Appendix D Action coverage of CAAP agent
Appendix E Additional related work
Vision models for agents
Numerous advanced techniques have been developed for detecting and interpreting UI elements on screens. Faster R-CNN (Ren et al., 2015) employs a Region Proposal Network (RPN) to extract areas of interest, followed by accurate object detection within these regions. This approach offers high accuracy in complex scenes, though at the cost of slower processing speed. The Single Shot MultiBox Detector (SSD) (Liu et al., 2016) offers rapid inference by eliminating the RPN and maintains decent accuracy. To improve small object detection and reduce false positives, one can diversify default boxes and feature maps for the SSD. Co-DETR (Zong et al., 2023), which leverages a Transformer-based architecture, performs detection while learning the relationships between objects. This anchor-free approach simplifies the learning process and effectively understands object interrelations, but it requires significant time and resources for training. YOLO (Redmon et al., 2016) processes predictions for object classes and bounding boxes across the entire image in one go. YOLO has both high speed and high accuracy compared to most other models. Unlike the previously mentioned technologies that primarily focus on extracting positional and typological information of objects, Pix2Struct (Lee et al., 2023) specializes in understanding detected objects. Designed around a Vision Transformer (ViT)(Dosovitskiy et al., 2020) and a T5 transformer(Raffel et al., 2020), it takes an image as input and outputs structured text, playing a crucial role in extracting and analyzing detailed attributes and embedded text information. Our approach leverages the combination of YOLO and Pix2Struct for visual information processing, aiming to capitalize on the strengths of both high-speed object detection and detailed attribute analysis.
Appendix F Case study of MiniWoB++ tasks
@@
Appendix G Annotation tool
@@
Appendix H Prompt examples
@@
Figure 7: A prompt example of generating a rationale for an agent action.
@@
### The UI Element List of the Current Screen That You Can Take Next Actions On ### (Note: Coordinates are given in the form: center_x [left_edge_x-right_edge_x], center_y [top_edge_y-bottm_edge_y]) - ID: element_1, data: {type: dropdown, X: 76 [1-152], Y: 66 [55-76], text: "Theodora", focused: False} - ID: element_2, data: {type: button, X: 50 [1-98], Y: 96 [80-112], text: "Submit", focused: False} ### Instructions ### What should be the next actions(action_2, action_3, …) that can be performed on the current screen? If there is an action that can complete the task, perform it immediately. It’s better if you can complete the task with fewer actions. Your answer must be composed of the following five sections: First, explain in detail what has been done so far up to action_1 and analyze why these steps were needed. Do not assume that the user could have made a mistake. Secondly, describe every single screen component that contains information that helps user solve the task or that needs to be interacted with. Explain about the components step by step in a detailed manner considering the given task. When the task deals with a list of items (e.g. finding the size of a group, identifying N-th item in order, etc.), you must include the full iteration of each and every items, like this: (1)first_item, (2)second_item, …, (N)N-th_item. Thirdly, describe what needs to be done to complete the task, detailing each action from start to finish. Then, identify which steps can currently be performed based on the UI elements visible on the screen, and mention the ID of these elements. If there is a demo available, first describe the sequence of actions demonstrated, then link these actions to the steps in your plan before outlining the complete action plan. Lastly, each action must be in the form of Action_2=(Action: functions.some_function_name, Argument: {property_name: property_val}). Return the actions that need to be performed on the current screen. The actions must be separated by new line characters. In case there are three actions to be performed, your response will be in the following form: Action_2=(Action: functions.some_function_name, Argument: {property_name: property_val}) Action_3=(Action: functions.some_function_name, Argument: {property_name: property_val}) Action_4=(Action: functions.some_function_name, Argument: {property_name: property_val}) - MAKE SURE that when you recommend actions that need to interact with UI elements, the UI elements MUST BE in the current screen. - MAKE SURE that when satisfying the task completion conditions, click the ’Submit’ button to finish. - MAKE SURE that even if you do not find a suitable action, return the most plausible one. - MAKE SURE that since texts in elements are extracted by recognizing with OCR, solve the given task considering that some mis-typos could exist.
Figure 8: An example of a CAAP prompt.
### Analysis of Action History So far, the only action taken has been the initiation of the task (action_1: {name: start}). This initial action is necessary to begin the process of selecting "Helli" from the dropdown list and submitting the choice. No specific interactions with the UI elements have been made yet. ### Description of Screen Components 1. **Dropdown Menu (element_1)**: This element is a dropdown menu currently displaying "Theodora" as the selected item. Its coordinates are given as X: 76 [1-152], Y: 66 [55-76], indicating its position on the screen. The dropdown is not focused, meaning it hasn’t been interacted with yet. This element is crucial for the task as it contains the list of names from which "Helli" needs to be selected. 2. **Submit Button (element_2)**: This is a button with the text "Submit" located at X: 50 [1-98], Y: 96 [80-112]. This button is essential for completing the task as it needs to be clicked after selecting "Helli" from the dropdown menu to confirm the selection. ### Completion Plan To complete the task, the following steps need to be taken: 1. Click on the dropdown menu to expand it and reveal the list of names. 2. Select "Helli" from the list of names in the dropdown. 3. Click the "Submit" button to confirm the selection. Based on the current screen, the actions that can be performed are: - Clicking on the dropdown menu (element_1) to expand it and reveal the list of names, including "Helli". - After selecting "Helli", clicking on the "Submit" button (element_2) will be necessary, but this can only be done after "Helli" has been successfully selected. Given the current screen elements, the next action is to interact with the dropdown menu to expand it and reveal the list of names. ### Actions to be Performed Action_2=(Action: functions.click_element, Argument: {element_id: 1})
@@
Appendix I Function calling example
@@
Appendix J Task-level result of MiniWoB++ Experiments
J.1 Per-task results for ablation study on CAAP components
Table 10: Per-task SRs with respect to varying combinations of prompt components.
@@
Task Full Demo without CoT phrases CoT phrases with No CoT,
configurations CoT phrases without demo action-only demo no demo
@@
find-greatest 1.000 1.000 0.420 0.860 1.000
@@
Average SR 0.945 0.922 0.850 0.910 0.812
J.2 Comparison of the CAAP and RCI promptings
Table 11: Per-task SRs for RCI agent with CAAP and original RCI agent.
Task RCI RCI
agent agent
with (original)
CAAP
@@
Task RCI RCI
agent agent
with (original)
CAAP
@@
Average SR 0.840 0.769
J.3 Comparison of the CAAP agent with SoTAs in per-task result
Table 12: Comparison with other works in terms of SR per task for MiniWoB++ problems.
@@
drag-circle 0.980 - - - - - - -
Task CAAP SeeClick Pix2Act $C$-Net WebGUM AdaPlanner RCI Human
@@
Task CAAP SeeClick Pix2Act $C$-Net WebGUM AdaPlanner RCI Human
@@
Average SR 0.945 0.694 0.962 0.936 0.925 0.929 0.940 0.935
Appendix K Task coverage of the agents that rely solely on screen images for MiniWoB++
Figure 11: Task coverage of the agents with image-only inputs for MiniWoB++